OpenStudy (anonymous):

Find the solution to the differential equation dy/dx+y/4=0 y=? Find the solution to the differential equation dy/dx+y/4=0 y=? @Mathematics

OpenStudy (amistre64):

i think you need on the them there e^(x/4)

OpenStudy (amistre64):

or just plain seperation of variables

OpenStudy (amistre64):

but the e is sooo cool

OpenStudy (anonymous):

This is a simple first-order homogenous equation. The characteristic equation of the solution is \[y = C_1 e^{-at}\] where a is the coefficient of the y term.

jimthompson5910 (jim_thompson5910):

dy/dx+y/4=0 dy/dx=-y/4 dy/y=-dx/4 Now integrate both sides and solve for y to get ln(|y|) = -x/4 |y| = e^(-x/4) y = +-e^(-x/4)

OpenStudy (amistre64):

yessss.... i knew it lol

OpenStudy (anonymous):

sorry wait

OpenStudy (amistre64):

dont forget that constant or itll mess you up

OpenStudy (anonymous):

Y(0)=12

OpenStudy (amistre64):

c = 12 then when x=0

OpenStudy (anonymous):

\[y = C_1 e^{-{t \over 4}}\] Set y=0 and t=0 and solve for C_1

OpenStudy (anonymous):

Sorry y=12 and t = 0

OpenStudy (anonymous):

no i want to do it jim thompsons way with Y(0)=12

OpenStudy (amistre64):

umm ... its all the same

OpenStudy (amistre64):

once you get to the y = Ce^-(x/4) you plug in and solve for C

OpenStudy (anonymous):

Dude e^0=1. 12 = c1.

OpenStudy (anonymous):

ok hold on, im trying to solve for C so give me a sec

OpenStudy (amistre64):

and i dont think y=+- e^.. is a solution to a differential equation

OpenStudy (amistre64):

something feels off there

OpenStudy (anonymous):

C=11 right?

OpenStudy (anonymous):

It could be but the sign depends on the sign of the coefficient. In this case, c is positive. It really should be written with a coefficient and not a sign.

jimthompson5910 (jim_thompson5910):

I got the +- from the fact that |y| = e^(-x/4)

OpenStudy (anonymous):

No. \[12 = C_1e^{-{0 \over 4}} = C_1e^0 = C_1\]

OpenStudy (amistre64):

from what i recall reading; in order to be a solution it has to be a function; and y=+- .... might be taboo in the rigidity of it, just a thought

jimthompson5910 (jim_thompson5910):

well that means y = e^(-x/4), y = -e^(-x/4)

OpenStudy (anonymous):

isnt it plus C? not multiplied by the e^(0/4)

OpenStudy (amistre64):

good luck ...

jimthompson5910 (jim_thompson5910):

dy/dx+y/4=0 dy/dx=-y/4 dy/y=-dx/4 Now integrate both sides and solve for y to get ln(|y|) = -x/4 + C |y| = e^(-x/4+C) y = e^(-x/4+C) or y = -e^(-x/4+C) y = e^(-x/4)*e^C or y = -e^(-x/4)*e^C y = e^C*e^(-x/4) or y = -e^C*e^(-x/4) y = k*e^(-x/4) or y = -k*e^(-x/4)

jimthompson5910 (jim_thompson5910):

Since y(0)=12, we know that the function y(x) is positive So y(x) > 0 which means that we pick the first function y = k*e^(-x/4) and plug in x = 0 and y = 12. From here solve for k y = k*e^(-x/4) 12 = k*e^(-0/4) 12 = k*e^(0) 12 = k*1 12 = k k = 12 So the function is y = 12*e^(-x/4)

OpenStudy (anonymous):

ya ok i got that

OpenStudy (anonymous):

thanks